Colony Inspection – 10th May

Observations:

Weather has started to warm up and temperatures are returning to normal – around 15C. Weather has been changeable with some heavy showers.

 

still lots of blossom – Chestnut now in full bloom, thyme, lilac

Hive Record
Colony
Q
BIAS
Brood
Stores
CommentNext Inspection
CyanYY8+8 still expandingAdded superGoing wellExchange manky frame at outside of brood box
DenimYY7 + DT - same as last inspectionStarting to work superHad emptied scraped frameRemove DT, re-mark queen
EmeraldYY8+7 - about same as previous inspectionSmall amount in SuperStarting to lay up DTCheck for space
FlaxNY7 frames of BIAS, so have expanded wellAdded superHave drawn 2 of the 4 frames foundationRemove DT
GreenYY5 frames BIAS, not expandingPlentyDT - 5:100 OKReplace manky frame and check for expansion
HoneydewNY6.5 biasStarting to work superBees a bit ratty, could be weatherFind and re-mark queen